# Finding all solutions of qKZ equations in characteristic $p$

### v1

@@ 7.1. Subbundle of qKZ connection
The $R$-matrix $R(u)\,=\,\frac{u-P}{u-1}$ has a pole if $u=1$ and is degenerate if $u=-1$. Denote by $H_{i,j,m}$ the affine hyperplane in ${\mathbb{K}}^{n}$ defined by the equation $z_{i}-z_{j}-m=0$ where $1\leqslant i<j\leqslant n$, $m\in{\mathbb{F}}_{p}$. Let $\bar{\mathcal{A}}^{\circ}$ be the arrangement in ${\mathbb{K}}^{n}$ of all hyperplanes $H_{i,j,m}$. Let $\bar{\mathcal{A}}={\mathbb{K}}^{n}-\bar{\mathcal{A}}^{\circ}$ denote its complement. The qKZ operators are well-defined over $\bar{\mathcal{A}}$ and are nondegenerate.
Assume that $10<d(\kappa)<n-$ 1. Then the $p$-hypergeometric solutions
(7.3) $Q^{\ell p-1}(z;\kappa),\qquad\ell=1,\dots,d(\kappa),$
define flat sections of the qKZ connection which we call the $p$-hypergeometric sections.
Recall the minors $M_{I}(z;\kappa)$ defined for any $I=\{1\leqslant i_{1}<\dots<i_{d(\kappa)}\leqslant n\}$ in Section 5.2 with the help of these $p$-hypergeometric sections. Denote by $\mathcal{A}(\kappa)$ the Zarisky open subset of $\bar{\mathcal{A}}$ consisting of points $b\in\bar{\mathcal{A}}$ such that at least one of the minors $M_{I}(z;\kappa)$ is nonzero at $b$.
@@
Notice that the minors $M_{I}(z;\kappa)$ are polynomials in $z$ with coefficients in ${\mathbb{F}}_{p}$ and are independent of the field ${\mathbb{K}}$. Notice also that the base $\mathcal{A}(\kappa)$ is invariant with respect to the affine translations,
$(z_{1},\dots,z_{n})\mapsto(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}),\qquad a=1,\dots,n.$
@@
If $d(\kappa)=0$, we define $\mathcal{A}(\kappa)=\bar{\mathcal{A}}$. In this case we define $\mathcal{S}(\kappa)\to\mathcal{A}(\kappa)$ to be the rank 0 subbundle of $V\times\mathcal{A}(\kappa)$ and also define $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ to be $V\times\mathcal{A}(\kappa)$.
7.2. Subbundle of dual qKZ connection
@@
Assume that $10<d(-\kappa)<n-$ 1. This assumption is equivalent to the assumption $10<d(\kappa)<n-$ 1. Consider the $p$-hypergeometric solutions $Q^{mp-1}(z;-\kappa)$, $m=1,\dots,d(-\kappa)$, of the qKZ equations step $-\kappa$. Then the $V$-valued polynomials
(7.4) $Q^{mp-1}(-z;-\kappa),\qquad m=1,\dots,d(-\kappa),$
define flat sections of the dual qKZ connection which we also call the $p$-hypergeometric sections.
Recall the minors $M_{I}(z;-\kappa)$ defined for any $I=\{1\leqslant i_{1}<\dots<i_{d(-\kappa)}\leqslant n\}$ in Section 5.2 with the help of the $p$-hypergeometric solutions $Q^{mp-1}(z;-\kappa)$, $m=1,\dots,d(-\kappa)$. Denote by $\mathcal{B}(\kappa)$ the Zarisky open subset of $\bar{\mathcal{A}}$ consisting of points $b\in\bar{\mathcal{A}}$ such that at least one of the polynomials $M_{I}(-z;-\kappa)$ is nonzero at $b$.
For any point $b\in\mathcal{B}(\kappa)$, the vectors $Q^{mp-1}(-b,-\kappa)$, $m=1,\dots,d(-\kappa)$, are linearly independent and span a $d(-\kappa)$-dimensional ${\mathbb{K}}$-vector subspace $\mathcal{S}^{*}(b,-\kappa)$ of $V$. These subspaces $\mathcal{S}^{*}(b,-\kappa)$, $b\in\mathcal{B}(\kappa)$, form a vector subbundle $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ of the trivial bundle $V\times\mathcal{B}(\kappa)\to\mathcal{B}(\kappa)$.
@@
(7.5) $S:\mathcal{S}^{*}(b,-\kappa)\otimes V/\mathcal{S}(b,\kappa)\to{\mathbb{K}}$
of the fibers of these bundles, by Theorem 6.1. The discrete connections on $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ and $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ are dual with respect to the Shapovalov form, that is, for any $u\in\mathcal{S}^{*}(b,-\kappa),$ $v\in V/\mathcal{S}(b,\kappa)$, and $a=1,\dots,n$, we have
$S(u,v)=S(K_{a}(-b;-\kappa)u\,,K_{a}(b\>;\kappa)v).$
@@
(7.6) $S(Q^{mp-1}(-b,-\kappa)\,,\,T^{\ell}(b,\kappa))=\delta_{\ell,m}\,,\qquad m=1,\dots,d(-\kappa).$
@@
Assume that $p>n$ and $10<d(-\kappa)\leqslant n-$ 1. Then $p$-quasi-hypergeometric sections $T^{\ell}(z;\kappa),$ $\ell=1,\dots,d(-\kappa)$, of the quotient bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$ form a flat basis of the space of sections of that bundle.
@@
(7.8) $T^{1}(z_{1},z_{2},2)=\left(\frac{3}{2z_{1}-2z_{2}+2}\,,\ \frac{3}{-2z_{1}+2z_{2}-2}\right).$
7.3. Reduced $p$-curvature operators
@@
If  $p>n$,  $\kappa\in{\mathbb{F}}_{p}^{\times}$,  and $10<d(\kappa)<n-$ 1, then the span of the $p$-hypergeometric sections $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, lies in the kernel of $\hat{C}_{a}(z,\kappa)$ and contains the image of $\hat{C}_{a}(z;\kappa)$ for every $a=1,\dots,n$.
If $d(\kappa)=p-1$ or  00, then all reduced $p$-curvature operators equal zero.
@@
We have
(7.9) $\hat{C}_{a}(z;\kappa)\hat{C}_{b}(z;\kappa)=0,\qquad a,b=1,\dots,n.$
Proof of Theorem 7.3. The span lies in the kernel of $\hat{C}_{a}(z;\kappa)$ by formula (2.8).
The operator $\hat{C}_{a}(z;\kappa)$ annihilates the span, hence $\hat{C}_{a}(z;\kappa)$ induces a well-defined operator on the fibers of the quotient bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$. This induced operator is the $a$-th reduced $p$-curvature operator of the qKZ connection on the quotient bundle. The quotient bundle has a flat basis of $p$-quasi-hypergeometric sections over the Zarisky open subset $\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$. Hence all reduced $p$-curvature operators of the qKZ connection on the quotient bundle are zero. Therefore the image of $\hat{C}_{a}(z;\kappa)$ is contained in the span.
@@
If  $p>n$,  $\kappa\in{\mathbb{F}}_{p}^{\times}$,  and $10<d(\kappa)<n-$ 1, then every reduced $p$-curvature operator $\tilde{C}_{a}(z;\kappa)$ is nonzero.
@@
Consider a normalized $p$-curvature operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ of the associated differential KZ equations. In a basis of $V$, the entries of the matrix of the operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ are homogeneous polynomials in $z$ of degree $(n-2)p$. By Corollary 2.2, in the same basis, the entries of the matrix of the operator $\tilde{C}_{a}(z,\kappa)$ are polynomials in $z$ of degree $(n-2)p$ whose top-degree parts equal the corresponding entries of the matrix of the operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$.
It is proved in [VV1, Theorem 1.13] that if $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$,  then every reduced $p$-curvature operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is a (nonzero) operator of rank 1. Hence every reduced $p$-curvature operator $\tilde{C}_{a}(z,\kappa)$ is a nonzero operator. ∎
@@
For $n=3$ we have $\dim V=2$. Let $p>3$ and $d(\kappa)=1$. Then $d(-\kappa)=1$. For $a=1,2,3$, the kernel of the reduced $p$-curvature operator $\hat{C}_{a}(z,\kappa)$ is generated by $Q^{p-1}(z,\kappa)$ and the image of $\hat{C}_{a}(z,\kappa)$ is generated by $Q^{p-1}(z,\kappa)$. Such an operator is determined uniquely up to multiplication by a scalar rational function in $z$.
For an operator $F:V\to V$, denote by $F^{*}:V\to V$ the operator dual to $F$ under the Shapovalov form, $S(F^{*}x,y)=S(x,Fy)$.
@@
(7.10) $\phantom{aaaaaa}\hat{C}_{a}(z,-\kappa)\,=\,-\,\hat{C}_{a}(-z;\kappa)^{*}.$
@@
We have
(7.11) $S(x,y)=S(C_{a}(-z;-\kappa)x,C_{a}(z;\kappa)y)$
by formulas (2.27) and (2.32). Hence
(7.12) $C_{a}(-z;-\kappa)=(C_{a}(z;\kappa)^{-1})^{*}.$
We also have $(C_{a}(z;\kappa)-1)^{2}=0$ by Corollary 7.4. Then
@@
(7.13) $\phantom{aaaaaa}\tilde{C}_{a}(z;-\kappa)$ $=$ $(-1)^{n}\tilde{C}_{a}(-z;\kappa)^{*}.$
@@
Formula (7.13) follows from equation (7.10) and the following formulas
@@
7.4. All solutions of qKZ equations for $\kappa\in{\mathbb{F}}_{p}^{\times}$
@@
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $10<d(\kappa)<n-$ 1. Let $f(z)$ be a $V$-valued rational function in $z$ which is a solution of the qKZ equations with step $\kappa$. Then $f(z)$ is a linear combination of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, with coefficients which are rational functions in $z_{i}^{p}-z_{i}$, $i=1,\dots,n$.
@@
For $a=1,\dots,n$, consider the normalized $p$-curvature operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ and $\tilde{C}_{a}(z,\kappa)$. Both of these operators are polynomials in $z$, and the polynomial $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is the top-degree part of the polynomial $\tilde{C}_{a}(z,\kappa)$. The polynomial $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is nonzero by [VV1, Theorem 1.13] and hence $\tilde{C}_{a}(z,\kappa)$ is a nonzero operator.
It was proved in [VV1, Theorem 1.8] that if $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $10<d(\kappa)<n-$ 1, then all solutions of the KZ equations are linear combinations of the $p$-hypergeometric solutions. Hence the intersection of kernels of the operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$, $a=1,\dots,n$, is of dimension $d(\kappa)$ for generic $z$, and the span of images of the operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$, $a=1,\dots,n$, is of dimension $n-1-d(\kappa)$ for generic $z$. Therefore, the span of images of the operators $\tilde{C}_{a}(z,\kappa)$, $a=1,\dots,n$, has dimension at least $n-1-d(\kappa)$ for generic $z$. This implies that the span of values of flat sections of the qKZ connection is of dimension not larger than $d(\kappa)$ for generic $z$. But we have $d(\kappa)$ flat linear independent $p$-hypergeometric sections $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$. Hence any flat section of the qKZ connection is a linear combination of the $p$-hypergeometric sections with 1-periodic coefficients. ∎
@@
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $10<d(\kappa)<n-$ 1. Let $f(z)$ and $g(z)$ be $V$-valued rational functions in $z$ where $f(z)$ is a solution the qKZ equations with step $\kappa$ and $g(z)$ is a solution the qKZ equations with step $-\kappa$. Then
(7.14) $S(g(-z),f(z))=0.$
@@
7.5. qKZ connection with $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$
@@
Let $p>n$ and $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$ Then all the normalized $p$-curvature operators $\tilde{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are nondegenerate for generic $z$.
@@
Formula (3.19) in [VV1] describes the spectrum of the $p$-curvature operators $C_{a}^{\operatorname{KZ}}(z;\kappa)$ of differential KZ equations. The formula shows that all $p$-curvature operators $C_{a}^{\operatorname{KZ}}(z;\kappa)$ are nondegenerate for generic $z$. In a basis of $V$, the matrices of $\tilde{C}_{a}^{\operatorname{KZ}}(z;\kappa)$ are homogeneous polynomials in $z$ of degree $(n-2)p$. Hence their determinants are nonzero homogeneous polynomials in $z$. By Corollary 2.2, the determinants of the normalized $p$-curvature operators $\tilde{C}_{a}(z;\kappa)$ are nonzero polynomials. The lemma follows. ∎

### v2

@@ 7.1 Subbundle of qKZ connection
The $R$-matrix $R(u)=\frac{u-P}{u-1}$ has a pole if $u=1$ and is degenerate if u=1u=-1. Denote by $H_{i,j,m}$ the affine hyperplane in ${\mathbb{K}}^{n}$ defined by the equation $z_{i}-z_{j}-m=0$ where $1\leqslant i<j\leqslant n$, $m\in{\mathbb{F}}_{p}$. Let $\bar{\mathcal{A}}^{\circ}$ be the arrangement in ${\mathbb{K}}^{n}$ of all hyperplanes $H_{i,j,m}$. Let $\bar{\mathcal{A}}={\mathbb{K}}^{n}-\bar{\mathcal{A}}^{\circ}$ denote its complement. The qKZ operators are well-defined over $\bar{\mathcal{A}}$ and are nondegenerate.
Assume that $0<d(\kappa)<n-1$. Then the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, define flat sections of the qKZ connection which we call the $p$-hypergeometric sections.
Recall the minors $M_{I}(z;\kappa)$ defined for any $I=\{1\leqslant i_{1}<\dots<i_{d(\kappa)}\leqslant n\}$ in Section 5.2 with the help of these $p$-hypergeometric sections. Denote by $\mathcal{A}(\kappa)$ the Zariski open subset of $\bar{\mathcal{A}}$ consisting of points $b\in\bar{\mathcal{A}}$ such that at least one of the minors $M_{I}(z;\kappa)$ is nonzero at $b$.
@@
Notice that the minors $M_{I}(z;\kappa)$ are polynomials in $z$ with coefficients in ${\mathbb{F}}_{p}$ and are independent of the field $\mathbb{K}$}. Notice also that the base $\mathcal{A}(\kappa)$ is invariant with respect to the affine translations, $(z_{1},\dots,z_{n})\mapsto(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})$, $a=1,\dots,n$.
@@
If $d(\kappa)=0$, we define $\mathcal{A}(\kappa)=\bar{\mathcal{A}}$. In this case, we define $\mathcal{S}(\kappa)\to\mathcal{A}(\kappa)$ to be the rank 0 subbundle of $V\times\mathcal{A}(\kappa)$ and also define $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ to be $V\times\mathcal{A}(\kappa)$.
7.2 Subbundle of dual qKZ connection
@@
Assume that $0<d(-\kappa)<n-1$. This assumption is equivalent to the assumption $0<d(\kappa)<n-1$. Consider the $p$-hypergeometric solutions $Q^{mp-1}(z;-\kappa)$, $m=1,\dots,d(-\kappa)$, of the qKZ equations step $\kappa$. Then the $V$-valued polynomials $Q^{mp-1}(-z;-\kappa)$, $m=1,\dots,d(-\kappa)$, define flat sections of the dual qKZ connection which we also call the $p$-hypergeometric sections.
Recall the minors $M_{I}(z;-\kappa)$ defined for any $I=\{1\leqslant i_{1}<\dots<i_{d(-\kappa)}\leqslant n\}$ in Section 5.2 with the help of the $p$-hypergeometric solutions $Q^{mp-1}(z;-\kappa)$, $m=1,\dots,d(-\kappa)$. Denote by $\mathcal{B}(\kappa)$ the Zariski open subset of $\bar{\mathcal{A}}$ consisting of points $b\in\bar{\mathcal{A}}$ such that at least one of the polynomials $M_{I}(-z;-\kappa)$ is nonzero at $b$.
For any point $b\in\mathcal{B}(\kappa)$, the vectors $Q^{mp-1}(-b,-\kappa)$, $m=1,\dots,d(-\kappa)$, are linearly independent and span a $d(-\kappa)$-dimensional $\mathbb{K}$}-vector subspace $\mathcal{S}^{*}(b,-\kappa)$ of space $V$. These subspaces $\mathcal{S}^{*}(b,-\kappa)$, $b\in\mathcal{B}(\kappa)$, form a vector subbundle $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ of the trivial bundle $V\times\mathcal{B}(\kappa)\to\mathcal{B}(\kappa)$}.
@@
$S\colon\ \mathcal{S}^{*}(b,-\kappa)\otimes V/\mathcal{S}(b,\kappa)\to{\mathbb{K}}$
of the fibers of these bundles, by Theorem 6.1. The discrete connections on $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ and $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$} are dual with respect to the Shapovalov form, that is, for any $u\in\mathcal{S}^{*}(b,-\kappa)$, $v\in V/\mathcal{S}(b,\kappa)$, and $a=1,\dots,n$, we have $S(u,v)=S(K_{a}(-b;-\kappa)u,K_{a}(b;\kappa)v)$.
@@
$S\big(Q^{mp-1}(-b,-\kappa),T^{\ell}(b,\kappa)\big)=\delta_{\ell,m},\qquad m=1,\dots,d(-\kappa).$
@@
Assume that $p>n$ and $0<d(-\kappa)\leqslant n-1$. Then $p$-quasi-hypergeometric sections $T^{\ell}(z;\kappa)$}, $\ell=1,\dots,d(-\kappa)$, of the quotient bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$ form a flat basis of the space of sections of that bundle.
@@
$T^{1}(z_{1},z_{2},2)=\left(\frac{3}{2z_{1}-2z_{2}+2},\frac{3}{-2z_{1}+2z_{2}-2}\right).$
7.3 Reduced $\boldsymbol{p}$-curvature operators
@@
If $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$, then the span of the $p$-hypergeometric sections $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, lies in the kernel of $\hat{C}_{a}(z,\kappa)$ and contains the image of $\hat{C}_{a}(z;\kappa)$ for every $a=1,\dots,n$.
If $d(\kappa)=p-1$ or 0, then all reduced $p$-curvature operators equal zero.
@@
We have $\hat{C}_{a}(z;\kappa)\hat{C}_{b}(z;\kappa)=0$, $a,b=1,\dots,n$.
Proof of Theorem 7.3..
The span lies in the kernel of $\hat{C}_{a}(z;\kappa)$ by formula (2.4).
The operator $\hat{C}_{a}(z;\kappa)$ annihilates the span, hence $\hat{C}_{a}(z;\kappa)$ induces a well-defined operator on the fibers of the quotient bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$. This induced operator is the $a$-th reduced $p$-curvature operator of the qKZ connection on the quotient bundle. The quotient bundle has a flat basis of $p$-quasi-hypergeometric sections over the Zariski open subset $\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$. Hence all reduced $p$-curvature operators of the qKZ connection on the quotient bundle are zero. Therefore the image of $\hat{C}_{a}(z;\kappa)$ is contained in the span.
@@
If $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$, then every reduced $p$-curvature operator $\tilde{C}_{a}(z;\kappa)$} is nonzero.
@@
Consider a normalized $p$-curvature operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ of the associated differential KZ equations. In a basis of $V$, the entries of the matrix of the operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ are homogeneous polynomials in $z$ of degree (n2)p(n-2)p. By Corollary 2.2, in the same basis, the entries of the matrix of the operator $\tilde{C}_{a}(z,\kappa)$ are polynomials in $z$ of degree (n2)p(n-2)p whose top-degree parts equal the corresponding entries of the matrix of the operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$.
It is proved in [8, Theorem 1.13] that if $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, then every reduced $p$-curvature operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is a (nonzero) operator of rank 1. Hence every reduced $p$-curvature operator $\tilde{C}_{a}(z,\kappa)$ is a nonzero operator. ∎
@@
For $n=3$, we have $\dim V=2$. Let $p>3$ and $d(\kappa)=1$. Then $d(-\kappa)=1$. For $a=1,2,3$, the kernel of the reduced $p$-curvature operator $\hat{C}_{a}(z,\kappa)$ is generated by $Q^{p-1}(z,\kappa)$ and the image of $\hat{C}_{a}(z,\kappa)$ is generated by $Q^{p-1}(z,\kappa)$. Such an operator is determined uniquely up to multiplication by a scalar rational function in $z$.
For an operator $F\colon V\to V$, denote by $F^{*}\colon V\to V$ the operator dual to $F$ under the Shapovalov form, $S(F^{*}x,y)=S(x,Fy)$.
@@
$\hat{C}_{a}(z,-\kappa)=-\hat{C}_{a}(-z;\kappa)^{*}.$ (7.1)
@@
We have $S(x,y)=S(C_{a}(-z;-\kappa)x,C_{a}(z;\kappa)y)$ by formulas (2.10) and (2.7). Hence $C_{a}(-z;-\kappa)=\big(C_{a}(z;\kappa)^{-1}\big)^{*}$. We also have $(C_{a}(z;\kappa)-1)^{2}=0$ by Corollary 7.4. Then
@@
$\tilde{C}_{a}(z;-\kappa)=(-1)^{n}\tilde{C}_{a}(-z;\kappa)^{*}.$ (7.2)
@@
Formula (7.2) follows from equation (7.1) and the following formulas:
@@
7.4 All solutions of qKZ equations for $\boldsymbol{\kappa\in{\mathbb{F}}_{p}^{\times}}$
@@
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$. Let $f(z)$ be a $V$-valued rational function in $z$ which is a solution of the qKZ equations with step $\kappa$. Then $f(z)$ is a linear combination of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, with coefficients which are rational functions in $z_{i}^{p}-z_{i}$, $i=1,\dots,n$.
@@
For $a=1,\dots,n$, consider the normalized $p$-curvature operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ and $\tilde{C}_{a}(z,\kappa)$. Both of these operators are polynomials in $z$, and the polynomial $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is the top-degree part of the polynomial $\tilde{C}_{a}(z,\kappa)$. The polynomial $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is nonzero by [8, Theorem 1.13] and hence $\tilde{C}_{a}(z,\kappa)$ is a nonzero operator.
It was proved in [8, Theorem 1.8] that if $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$, then all solutions of the KZ equations are linear combinations of the $p$-hypergeometric solutions. Hence the intersection of kernels of the operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$, $a=1,\dots,n$, is of dimension $d(\kappa)$ for generic $z$, and the span of images of the operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$, $a=1,\dots,n$, is of dimension $n-1-d(\kappa)$} for generic $z$. Therefore, the span of images of the operators $\tilde{C}_{a}(z,\kappa)$, $a=1,\dots,n$, has dimension at least $n-1-d(\kappa)$ for generic $z$. This implies that the span of values of flat sections of the qKZ connection is of dimension not larger than $d(\kappa)$ for generic $z$. But we have $d(\kappa)$ flat linear independent $p$-hypergeometric sections $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$. Hence any flat section of the qKZ connection is a linear combination of the $p$-hypergeometric sections with 1-periodic coefficients. ∎
@@
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$. Let $f(z)$ and $g(z)$ be $V$-valued rational functions in $z$ where $f(z)$ is a solution the qKZ equations with step $\kappa$ and $g(z)$ is a solution the qKZ equations with step $\kappa$. Then
$S(g(-z),f(z))=0.$ (7.3)
@@
7.5 qKZ connection with $\boldsymbol{\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}}$
@@
Let $p>n$ and $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$ . Then all the normalized $p$-curvature operators $\tilde{C}_{a}(z;\kappa)$}, $a=1,\dots,n$, are nondegenerate for generic $z$.
@@
Formula (3.19) in [8] describes the spectrum of the $p$-curvature operators $C_{a}^{\operatorname{KZ}}(z;\kappa)$ of differential KZ equations. The formula shows that all $p$-curvature operators $C_{a}^{\operatorname{KZ}}(z;\kappa)$ are nondegenerate for generic $z$. In a basis of $V$, the matrices of $\tilde{C}_{a}^{\operatorname{KZ}}(z;\kappa)$ are homogeneous polynomials in $z$ of degree (n2)p(n-2)p. Hence their determinants are nonzero homogeneous polynomials in $z$. By Corollary 2.2, the determinants of the normalized $p$-curvature operators $\tilde{C}_{a}(z;\kappa)$ are nonzero polynomials. The lemma follows. ∎
@@
Acknowledgements
E.M. was supported in part by Simons Foundation grant 709444. A.V. was supported in part by Simons Foundation grant TSM-00012774. The authors thank P. Etingof, R. Rimanyi, and V. Vologodsky for useful discussions.